OpenStudy (some.random.cool.kid):

Circle A has center of (2, 3), and a radius of 5 and circle B has a center of (1, 4), and a radius of 10. What steps will help show that circle A is similar to circle B? Dilate circle A by a scale factor of 2. Translate circle A using the rule (x + 1, y − 1). Rotate circle A 180° about the center. Reflect circle A over the y-axis.

OpenStudy (some.random.cool.kid):

any help?

OpenStudy (some.random.cool.kid):

Havent used this site in months figured i would give it a whirl.

OpenStudy (some.random.cool.kid):

@jim_thompson5910 @zepdrix

OpenStudy (some.random.cool.kid):

@SolomonZelman @sleepyjess

jimthompson5910 (jim_thompson5910):

If you dilate figure A and you're able to match it to figure B (after dilation) then that shows figure A is similar to figure B For example, these two triangles are similar. The large one has all sides multiplied by the same constant to get the triangle bigger. http://mathworld.wolfram.com/images/eps-gif/Dilation_900.gif

OpenStudy (some.random.cool.kid):

well i put A? would i be correct?

OpenStudy (some.random.cool.kid):

Especially since you summed it up that way? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

yes A is correct

OpenStudy (some.random.cool.kid):

thanks, I may need one more.

OpenStudy (some.random.cool.kid):

A circle is centered at (7, 8) and has a radius of 11. Which of the following is the equation for this circle? (x − 7)2 + (y − 8)2 = 121 (x − 7)2 + (y − 8)2 = 11 (x + 7)2 + (y + 8)2 = 121 (x + 7)2 + (y + 8)2 = 11 I put c @jim_thompson5910 I am not sure this is correct though honestly becuase none of my research matches this equation. I think I have a broad idea but I cant pin point one.

OpenStudy (some.random.cool.kid):

@sweetburger @SamsungFanBoy

OpenStudy (some.random.cool.kid):

@jabez177

jimthompson5910 (jim_thompson5910):

`A circle is centered at (7, 8)` h = 7 k = 8 (x-h)^2 + (y-k)^2 = r^2 notice how it's x-h and y-k. The negatives in front of h and k tell us to take the opposite (x-7)^2 + (y-8)^2 = r^2 then you'd plug in the given radius for r

OpenStudy (some.random.cool.kid):

was 7 and 8 the given radius?

jimthompson5910 (jim_thompson5910):

no, the radius is given after the center point is given

jimthompson5910 (jim_thompson5910):

(7,8) is your center

OpenStudy (some.random.cool.kid):

So we are trying to find the radius? They didn't tell us? If so it could be 11.

jimthompson5910 (jim_thompson5910):

yeah they just gave it to you. The radius is 11

jimthompson5910 (jim_thompson5910):

if r = 11, then r^2 = ???

OpenStudy (some.random.cool.kid):

121 :)

OpenStudy (some.random.cool.kid):

so was c correct?

jimthompson5910 (jim_thompson5910):

go back to my first post

jimthompson5910 (jim_thompson5910):

the one where I talk about (x-h)^2+(y-k)^2

OpenStudy (some.random.cool.kid):

(x − 7)2 + (y − 8)2 = 121 ^ ... was I supposed to change the opposite to this... v (x + 7)2 + (y + 8)2 = 121 Or the other way around thats where I am confused.

OpenStudy (some.random.cool.kid):

I thought based on your negtives it would be positives... but now your saying no i think? Idk

jimthompson5910 (jim_thompson5910):

Given center is (h,k) = (7,8). Radius is r = 11. So, \[\Large {\color{blue}{h = 7}}\] \[\Large {\color{green}{k = 8}}\] \[\Large {\color{red}{r = 11}}\] Plug them into the general circle equation \[\Large (x-h)^2 + (y-k)^2 = r^2\] \[\Large (x-{\color{blue}{h}})^2 + (y-{\color{green}{k}})^2 = ({\color{red}{r}})^2\] \[\Large (x-{\color{blue}{7}})^2 + (y-{\color{green}{8}})^2 = ({\color{red}{11}})^2\] \[\Large (x-7)^2 + (y-8)^2 = 121\]

OpenStudy (some.random.cool.kid):

oh so it is the negatives ok?

OpenStudy (some.random.cool.kid):

(x − 7)2 + (y − 8)2 = 121

OpenStudy (some.random.cool.kid):

^ like that?

jimthompson5910 (jim_thompson5910):

yes but don't say (x-7)2 write (x-7)^2

jimthompson5910 (jim_thompson5910):

`^` means exponent

OpenStudy (some.random.cool.kid):

oh i was using that as an arrow.

OpenStudy (some.random.cool.kid):

mb

jimthompson5910 (jim_thompson5910):

example 3x^2 = \(\Large 3x^2\)